what is the force required for a steel wire to double it length when it's area cross-section is 1 square cm and Young modulus is 200GPa as the length of the wire is double the change in length is equal to the original length.


1
Expert's answer
2021-10-20T10:42:45-0400

Determine strain:


"\epsilon=\frac{\Delta L}L=\frac{\sigma}E=1,\\\space\\
\sigma=\frac FA,\\\space\\
F=\sigma A=EA=20·10^6\text{ N}."